OpenStudy (anonymous):

The function f(x) = −x^2 + 44x − 384 models the profit, in dollars, a game developer makes for selling a new game, where x is the number of games sold and f(x) is the amount of profit. Part A: Determine the vertex. What does this calculation mean in the context of the problem? (5 points) Part B: Determine the x-intercepts. What do these values mean in the context of the problem? (5 points

OpenStudy (whpalmer4):

This is a parabola: \[y = ax^2 + bx + c\] The vertex of a parabola written in that form can be found to be \[x = -\frac{b}{2a},\ y = f(x)\]in other words, find the \(x\) value from the first formula, and plug it into the equation to get the \(y\) value there. The \(x\)-intercepts are the points where the parabola crosses the \(x\)-axis. This is the same as solving for \(x\) in \[f(x) = 0\]or\[-x^2+44x-384=0\]